OpenStudy (naomeeprz93):

96/20 into a mixed number?

OpenStudy (mieshapitts):

4 4/5 ?

OpenStudy (tkhunny):

Have you considered division?

OpenStudy (greenrune113):

96/20 How many times does 20 go into 96? 4 times - which equal to 80 96 - 80 = 16 \[4 \frac{ 16 }{ 20 }\] Simplify 16 over 20 by diving each by 4 16/4 = 4 20/4 5 \[4\frac{ 4 }{ 5 }\]